# Bridging Deep Reinforcement Learning and Motion Planning for Model-Free Navigation in Cluttered Environments

### v2

@@ Abstract
Deep Reinforcement Learning (DRL) has emerged as a powerful model-free paradigm for learning optimal policies. However, in navigation tasks with cluttered environments, DRL methods often suffer from insufficient exploration, especially under sparse rewards or complex dynamics with system disturbances. To address this challenge, we bridge general graph-based motion planning and DRL, enabling agents to explore cluttered spaces more effectively and achieve desired navigation performance. Specifically, we design a dense reward function grounded in a graph structure that spans the entire state space. This graph provides rich guidance, steering the agent toward optimal strategies. We validate our approach in challenging environments, demonstrating substantial improvements in exploration efficiency and task success rates.
@@ 1 Introduction
Deep Reinforcement Learning (DRL) provides a flexible and general approach for acquiring control policies in tasks characterized by high-dimensional observations and complex, uncertain dynamics [1, 2]. Its model-free nature enables agents to learn directly from interactions, eliminating the need for explicit dynamics modeling. This makes DRL particularly suitable for domains such as robotic manipulation [3, 4], autonomous driving [5], and locomotion [6], where accurate models can be difficult to obtain. In these applications, DRL demonstrates strong potential for generalization and robustness, supporting intelligent decision-making under diverse and dynamic conditions.
However, the practical success of DRL remains severely limited by its sample inefficiency, primarily due to inadequate exploration in sparse-reward or high-dimensional settings [7, 8]. Without informative feedback, agents may spend substantial time interacting with irrelevant states. Numerous strategies have been proposed to improve exploration in DRL, including intrinsic rewards [9, 10], density-based bonuses [11, 8], entropy regularization [12], policy perturbation [13], and hierarchical or logic-guided approaches [14, 15]. These approaches have partially alleviated exploration challenges and improved sample efficiency in regular environments. However, these methods still struggle with sparse reward signals and high-dimensional, cluttered environments—challenges that remain largely unsolved.
@@ 3 Preliminary and Problem Formulation
System Disturbance: We consider a continuous-time dynamical system $S$ with general unknown dynamics, whose evolution is described by: $\frac{ds}{dt}=\dot{s}=f(s,a,d)$ where $s\in\mathcal{S}\subseteq\mathbb{R}^{n}$ represents the state variables, $a\in\mathcal{A}\subseteq\mathbb{R}^{m}$ denotes the control inputs, and $d$ is the disturbance function.
@@ 4.2 Motion Planning Guided Reward
We adopt the definition of history from [26] with a slight modification $h_{t}=\{s_{1},a_{1},\dots,s_{t-1},a_{t-1}\}$ instead of $\{s_{1},a_{1},\dots,s_{t-1},a_{t-1},s_{t}\}$ , thereby the $D_{min}$ over a state-action sequence $\tau_{t}=h_{t}\bigcup\{s_{t},a_{t}\}$ is defined as $D_{min}(\mathbf{\tau})=\min_{s\in\mathbf{s}}{D(s)}$, where $\mathbf{s}$ denotes the state sequence in trajectory $\tau_{t}$. Ideally, we would check the change of $D_{min}$ at each step to determine whether the agent has moved closer to the goal region, i.e., check if $D_{min}(\tau_{t})<D_{min}(\tau_{t-1})$, and assign a positive reward $R_{+}$ accordingly. However, this design introduces a non-Markovian reward issue. In other words, $r(h_{t},s_{t},a_{t})\neq r(h^{\prime}_{t},s_{t},a_{t})$ could happen with different history $h_{t}$ and $h^{\prime}_{t}$. To address this problem, we denote $\mathcal{S}^{\times}=\{(s,D_{min})\mid s\in\mathcal{S},\,D_{min}\in\mathbb{R}\}$, so that $D_{min}$ can be directly retrieved from the augmented state. Correspondingly, the augmented reward function is defined as $\tilde{r}:\mathcal{S}^{\times}\times\mathcal{A}\times\mathcal{S}^{\times}\to\mathbb{R}$,
@@
$\theta^{*}=\arg\min_{\theta}\mathbb{E}_{(s^{\times},a,r,s^{\prime\times})\sim\mathcal{D}}\left[\left(Q_{\omega}(s,a)-y\right)^{2}\right],$
@@ 4.4 Advantages and Extensions
With the graph-based waypoint covering the whole environment, we could naturally extend the training objective from single start point to arbitrary start points. This brings some desirable properties that enable multiple extensions.
@@ 5 Experimental Results
Efficient Exploration: We consider each training task as a compositional reinforcement learning problem, where the agent must learn reusable primitives that enable reaching a goal from any designated region within the environment. Training is deemed successful once the agent reaches the goal region from all ten predefined test starting positions. To evaluate learning efficiency, we define the consumption as the global step at which this success criterion is first achieved and the total training budget. A lower consumption indicates more efficient exploration. If the agent fails to meet the success condition within the allotted training steps, a consumption value of highest global steps is assigned. As shown in Fig. 3 and Fig. 4, although the agent without guidance is occasionally able to reach the goal from certain starting positions, nearly all of them fail to satisfy the overall task criterion, which is reflected in their exceedingly high step consumption.
@@
Generalization to Unseen Situations: As discussed in Sec. 4.4, a key advantage of our framework lies in its ability to generalize to arbitrary start states and integrate recovery mechanisms seamlessly. To evaluate this capability, we conduct a visualization experiment as shown in Fig. 5. The left panel displays the discounted success score, defined as $\mathbb{I}_{\text{success}}\cdot(\text{steps})^{\gamma}$, where $\mathbb{I}_{\text{success}}$ is the indicator function for goal-reaching and $\gamma\in(0,1)$ is a discount factor that emphasizes earlier success. This illustrates that the agent learns a globally coherent value landscape that supports goal-reaching from diverse initializations. In the middle panel, we visualize 10 representative trajectories from randomly selected start points, all successfully reaching the goal without any external intervention. To simulate realistic failures or unexpected conditions, we introduce unseen obstacles in the test environment and activate a simple distance-based safety filter to avoid entering unsafe zones. The resulting trajectories, shown in the right panel, demonstrate that the agent can dynamically adapt its path and still reach the goal, confirming the effectiveness of our method in handling perturbed and previously unseen situations.
@@ Appendix A Cost-to-Go Calculation for Planning-Based Rewards
Discretization
The function DiscretizeEnv divides the continuous environment $\mathcal{X}$ into a uniform grid. Each cell is classified as either free or occupied, depending on whether it overlaps with an obstacle. The resulting map structure serves as the input graph for A* search.
Neighbors
Given a state $s$ and a map structure Map, the function Neighbors$(s,\text{Map})$ returns the set of neighboring states that are directly reachable from $s$. In a 4-connected grid, these are the adjacent horizontal and vertical cells; in an 8-connected grid, diagonal neighbors are also included. In a graph like RRG, neighbors are determined by edges in $E$.
Collision check
The function isObstacle$(s)$ returns true if the state $s$ is within an obstacle region, and false otherwise. This function is used to prune invalid neighbors from expansion.
Cost
The function Cost$(s,s^{\prime})$ returns the traversal cost between two adjacent states $s$ and $s^{\prime}$. This is typically set to the Euclidean or Manhattan distance between the states for grid maps, or to edge weights in roadmap graphs.
Heuristic
The function $h(s)$ provides an admissible estimate of the cost from state $s$ to the goal state $s_{\text{goal}}$. Common choices include Manhattan or Euclidean distances in grid settings.
Path reconstruction
Once the goal state is reached, the function ReconstructPath$(s_{\text{goal}})$ traces backward through the parent pointers to construct the shortest path from $s_{\text{start}}$ to $s_{\text{goal}}$. The result is a trajectory $\mathbf{X}^{*}=\{s_{0},s_{1},\dots,s_{d}\}$ such that $s_{0}=s_{\text{start}}$ and $s_{d}=s_{\text{goal}}$.
@@ Appendix B Theorem 1 proof
We first prove this reward is non-Markovian in the original state space $\mathcal{S}$. With the definition of MDP [2], we prove it by a contradiction. Assume, for the sake of contradiction, that the reward function in the original state space is Markovian. That is, at each timestep $t$
@@ Appendix C Proposition 1 Proof
We first show that there exists an optimal policy that reaches the goal under the original reward $\hat{r}$. For the MDP we described in Section 3, the existence of such a policy is guaranteed by standard results in measure-theoretic Markov decision processes (see, e.g., Theorem 3.2 in [34] and Theorem 8.7.2 in [26]).
@@
In sparse-reward settings, most transitions yield $\hat{r}=0$, and thus the shaping term dominates the learning signal. If $\Phi$ is positively correlated with goal proximity, then $\delta^{\prime}_{t}$ reflects meaningful progress even when the environment provides no extrinsic reward. According to the convergence theory of TD learning [2], larger TD errors, which remain proper learning rates, lead to faster value function updates. Therefore, our augmented rewards enhance early-stage TD signals and accelerate convergence in sparse environments.

### v3

@@ Abstract
Deep Reinforcement Learning (DRL) has emerged as a powerful model-free paradigm for learning optimal policies. However, in navigation tasks with cluttered environments, DRL methods often suffer from insufficient exploration, especially under sparse rewards or complex dynamics with system disturbances. To address this challenge, we bridge general graph-based motion planning and DRL, enabling agents to explore cluttered spaces more effectively and achieve desired navigation performance. Specifically, we design a dense reward function grounded in a graph structure that spans the entire state space. This graph provides rich guidance, steering the agent toward optimal strategies. We validate our approach in challenging environments, demonstrating substantial improvements in exploration efficiency and task success rates. The project website is available at: https://plen1lune.github.io/overcome_exploration/
@@ 1 Introduction
Deep Reinforcement Learning (DRL) provides a flexible and general approach for acquiring control policies in tasks characterized by high-dimensional observations and complex, uncertain dynamics [1, 2]. Its model-free nature enables agents to learn directly from interactions, eliminating the need for explicit dynamics modeling. This makes DRL particularly suitable for domains such as robotic manipulation [3, 4], autonomous driving [5], and locomotion [6], where accurate models can be difficult to obtain. In recent studies, DRL demonstrates strong potential for generalization and robustness, supporting intelligent decision-making under diverse and dynamic conditions [7].
However, the practical success of DRL remains severely limited by its sample inefficiency, primarily due to inadequate exploration in sparse-reward or high-dimensional settings [8, 9, 10]. Without informative feedback, agents may spend substantial time interacting with irrelevant states. Numerous strategies have been proposed to improve exploration in DRL, including intrinsic rewards [11, 12], density-based bonuses [13, 9], entropy regularization [14], policy perturbation [15], and hierarchical or logic-guided approaches [16, 17]. These approaches have partially alleviated exploration challenges and improved sample efficiency in regular environments. However, these methods still struggle with sparse reward signals and high-dimensional, cluttered environments—challenges that remain largely unsolved.
@@ 3 Preliminary and Problem Formulation
System Disturbance: We consider a continuous-time dynamical system $S$ with general unknown dynamics, whose evolution is described by $\frac{ds}{dt}=\dot{s}=f(s,a,d)$ where $s\in\mathcal{S}\subseteq\mathbb{R}^{n}$ represents the state variables, $a\in\mathcal{A}\subseteq\mathbb{R}^{m}$ denotes the control inputs, and $d$ is the disturbance function.
@@ 4.2 Motion Planning Guided Reward
We adopt the definition of history from [28] with a slight modification $h_{t}=\{s_{1},a_{1},\dots,s_{t-1},a_{t-1}\}$ instead of $\{s_{1},a_{1},\dots,s_{t-1},a_{t-1},s_{t}\}$ , thereby the $D_{min}$ over a state-action sequence $\tau_{t}=\{h_{t},s_{t},a_{t}\}$ is defined as $D_{min}(\mathbf{\tau})=\min_{s\in\mathbf{s}}{D(s)}$, where $\mathbf{s}$ denotes the state sequence in trajectory $\tau_{t}$. Ideally, we would check the change of $D_{min}$ at each step to determine whether the agent has moved closer to the goal region, i.e., check if $D_{min}(\tau_{t})<D_{min}(\tau_{t-1})$, and assign a positive reward $R_{+}$ accordingly. However, this design introduces a non-Markovian reward issue. In other words, $r(h_{t},s_{t},a_{t})\neq r(h^{\prime}_{t},s_{t},a_{t})$ could happen with different history $h_{t}$ and $h^{\prime}_{t}$. To address this problem, we denote $\mathcal{S}^{\times}=\{(s,D_{min})\mid s\in\mathcal{S},\,D_{min}\in\mathbb{R}\}$, so that $D_{min}$ can be directly retrieved from the augmented state. Correspondingly, the augmented reward function is defined as $\tilde{r}:\mathcal{S}^{\times}\times\mathcal{A}\times\mathcal{S}^{\times}\to\mathbb{R}$,
@@
$P\bigl{(}\tilde{r}_{t}\mid s_{0}^{\times},a_{0},\dots,s_{t}^{\times},a_{t},s_{t+1}^{\times}\bigr{)}\;=\;P\bigl{(}\tilde{r}_{t}\mid s_{t}^{\times},a_{t},s_{t+1}^{\times}\bigr{)}$
@@
$\theta^{*}=\arg\min_{\theta}\mathbb{E}_{(s^{\times},a,r,s^{\prime\times})\sim\mathcal{D}}\left[\left(Q_{\omega}(s,a)-y\right)^{2}\right]$
@@ 4.4 Advantages and Extensions
With the graph-based waypoints covering the whole environment, we could naturally extend the training objective from single start point to arbitrary start points. This brings some desirable properties that enable multiple extensions.
@@ 5 Experimental Results
Efficient Exploration: We consider each training task as a compositional reinforcement learning problem, where the agent must learn reusable primitives that enable reaching a goal from any designated region within the environment. Training is deemed successful once the agent reaches the goal region from all ten predefined test starting positions. To evaluate learning efficiency, we define consumption as the global step at which this success criterion is first achieved and the total training budget. A lower consumption indicates more efficient exploration. If the agent fails to meet the success condition within the allotted training steps, a consumption value of highest global steps is assigned. As shown in Fig. 3 and Fig. 4, although the agent without guidance is occasionally able to reach the goal from certain starting positions, nearly all of them fail to satisfy the overall task criterion, which is reflected in their exceedingly high step consumption.
@@
Generalization to Unseen Situations: As discussed in Sec. 4.4, a key advantage of our framework lies in its ability to generalize to arbitrary start states and integrate recovery mechanisms seamlessly. To evaluate this capability, we conduct a visualization experiment as shown in Fig. 5. The left panel displays the discounted success score, defined as $\mathbb{I}_{\text{success}}\cdot(\text{steps})^{\gamma}$, where $\mathbb{I}_{\text{success}}$ is the indicator function for goal-reaching, “steps” denotes the episode length and $\gamma\in(0,1)$ is a discount factor that emphasizes earlier success. This illustrates that the agent learns a globally coherent value landscape that supports goal-reaching from diverse initializations. In the middle panel, we visualize 10 representative trajectories from randomly selected start points, all successfully reaching the goal without any external intervention. To simulate realistic failures or unexpected conditions, we introduce unseen obstacles in the test environment and activate a simple distance-based safety filter to avoid entering unsafe zones. The resulting trajectories, shown in the right panel, demonstrate that the agent can dynamically adapt its path and still reach the goal, confirming the effectiveness of our method in handling perturbed and previously unseen situations. While not the primary focus, this illustrates the potential of our framework for supporting broader safety-critical extensions.
@@ 6 Conclusion
Limitations and Future Work
Our method relies on a precomputed roadmap (e.g., RRG) to provide topological guidance. Although the roadmap does not need to be globally optimal, its connectivity can influence path diversity and task coverage, particularly in extremely sparse or disconnected environments. In practice, we assume that the environment is connected and continue sampling waypoints until sufficient connectivity is achieved. Although this heuristic may not guarantee formal completeness, we find that it works reliably in all our tested environments and does not significantly impact the learning or success rate in practice.
Another experiment limitation lies in the demonstration of simple recovery mechanisms, such as distance-based filters, which may not capture more nuanced safety constraints in highly dynamic or safety-critical domains. Nonetheless, our framework is compatible with more advanced model-free or model-based recovery modules, which we leave as promising future extensions.
While our method generalizes prior work from fixed start-goal pairs to arbitrary start positions, fully randomized goal configurations remain an open challenge. For each new goal, the guidance graph must be re-searched or re-planned, and storing separate graphs for all goals is impractical. We are investigating scalable alternatives and will pursue this in future work. Despite these limitations, we find that our method demonstrates strong progress in learning temporal-logic-specified primitives, and view LTL-guided generalization as a promising direction for future work.
We acknowledge the absence of real-world experiments in this work but our method is designed with practicality in mind and does not rely on simulation-specific assumptions. We need to emphasize that the proposed framework is an exploration enhancement strategy that is fundamentally model-agnostic. It can be naturally integrated into a wide range of RL algorithms, regardless of the underlying dynamics. In future work, we plan to apply this framework to more task-specific scenarios and include real-world deployment as part of our experimental agenda. Additionally, we do not compare with model-based approaches, as our focus is on model-free settings where environment dynamics are unknown or difficult to estimate.
@@ Appendix A Cost-to-Go Calculation for Planning-Based Rewards
Discretization The function DiscretizeEnv divides the continuous environment $\mathcal{X}$ into a uniform grid. Each cell is classified as either free or occupied, depending on whether it overlaps with an obstacle. The resulting map structure serves as the input graph for A* search.
Neighbors Given a state $s$ and a map structure Map, the function Neighbors$(s,\text{Map})$ returns the set of neighboring states that are directly reachable from $s$. In a 4-connected grid, these are the adjacent horizontal and vertical cells; in an 8-connected grid, diagonal neighbors are also included. In a graph like RRG, neighbors are determined by edges in $E$.
Collision check The function isObstacle$(s)$ returns true if the state $s$ is within an obstacle region, and false otherwise. This function is used to prune invalid neighbors from expansion.
Cost The function Cost$(s,s^{\prime})$ returns the traversal cost between two adjacent states $s$ and $s^{\prime}$. This is typically set to the Euclidean or Manhattan distance between the states for grid maps, or to edge weights in roadmap graphs.
Heuristic The function $h(s)$ provides an admissible estimate of the cost from state $s$ to the goal state $s_{\text{goal}}$. Common choices include Manhattan or Euclidean distances in grid settings.
Path reconstruction Once the goal state is reached, the function ReconstructPath$(s_{\text{goal}})$ traces backward through the parent pointers to construct the shortest path from $s_{\text{start}}$ to $s_{\text{goal}}$. The result is a trajectory $\mathbf{X}^{*}=\{s_{0},s_{1},\dots,s_{d}\}$ such that $s_{0}=s_{\text{start}}$ and $s_{d}=s_{\text{goal}}$.
@@ Appendix B Theorem 1 proof
We first prove this reward is non-Markovian in the original state space $\mathcal{S}$. With the definition of MDP [2], we prove it by a contradiction.
Let the full history up to $t-1$ be $h_{t-1}:=\{s_{0},a_{0},\dots,s_{t-1},a_{t-1}\}$. Assume, for the sake of contradiction, that the reward function in the original state space is Markovian. That is, at each timestep $t$
@@ Appendix C Proposition 1 Proof
We first show that there exists an optimal policy that reaches the goal under the original reward $\hat{r}$. For the MDP we described in Sec. 3, the existence of such a policy is guaranteed by standard results in measure-theoretic Markov decision processes (see, e.g., Theorem 3.2 in [36] and Theorem 8.7.2 in [28]).
@@
In sparse-reward domains $\hat{r}_{t}=0$ for most transitions, so the shaping term $F_{t}$ dominates $\delta^{\prime}_{t}$. Because $F_{t}>0$ precisely when the agent reduces its distance to the goal, the TD error now carries informative progress signals at every step. Provided that the learning-rate schedule $\{\alpha_{t}\}$ satisfies the usual Robbins–Monro conditions, the larger magnitude of these early TD errors increases the expected parameter update norm and empirically accelerates convergence, a phenomenon consistent with the convergence analysis of TD learning in Sutton and Barto [2]. ∎
Appendix D Experimental Setup
We evaluate three widely-used continuous-control algorithms—PPO, SAC, and TD3-on two custom environments: a Quadrotor navigation task (3-DoF velocity commands) and a 2-D Dog/Car navigation task (forward velocity and yaw rate). Figure 6 depicts both environments.
(a) Quadrotor navigation environment.
(b) Dog/Car-2D maze environment.
Figure 6: (a): a four-level $1\,\text{m}^{3}$ cubic maze for the Quadrotor agent, which navigates with 3-DoF velocity commands $\langle v_{x},v_{y},v_{z}\rangle$. (b): a $12\times 12$ planar maze with additional loops for the Dog/Car-2D agent, controlled by forward speed and yaw rate $\langle v,\omega\rangle$. In both tasks the agent starts from a start region and must reach a designated goal without collisions.
The quadrotor operates in a maze-like workspace, which is a $1\,\mathrm{m}\times 1\,\mathrm{m}\times 1\,\mathrm{m}$ cube discretised into four horizontal floors, each $0.25$m high. Every floor is a $4\times 4$ room grid ($0.25$m per room). A single vertical shaft (highlighted in green at construction time) links one designated room on each floor, enabling upward traversal while all other ceiling tiles remain closed. Actions are high-level commands $\langle v_{x},v_{y},v_{z}\rangle\in[-1.5,1.5]^{3}$m/s.
The ground agent navigates in a $12\times 12$ perfect maze augmented with $15\%$ additional loops (wall thickness $0.02$m, corridor width $0.14$m). Actions are high-level commands $\langle v,\omega\rangle$ with $v\in[0,1.0]$m/s and $\omega\in[-2.0,2.0]$rad/s. Episodes start in the bottom-left green square ($0.32$m ×  $0.32$m) and end in the top-right orange square of the same size.
Quadrotor agents are trained for $8{\times}10^{6}$ environment steps, while Dog/Car-2D agents are trained for $4{\times}10^{6}$ steps. All runs use four synchronous workers and log to Weights  Biases. Tables 2 and 3 list the remaining hyper-parameters, which are shared across all baseline algorithms within each environment to ensure fair comparison.
Table 2: Hyper-parameters for the Quadrotor environment.
Hyper-parameter PPO SAC TD3
Policy network [64, 64]6464[64,\,64][ 64 , 64 ] [64, 64]6464[64,\,64][ 64 , 64 ] [64, 64]6464[64,\,64][ 64 , 64 ]
Value / $Q$ network [64, 64]6464[64,\,64][ 64 , 64 ] [64, 64]6464[64,\,64][ 64 , 64 ] [64, 64]6464[64,\,64][ 64 , 64 ]
Learning rate $3{\times}10^{-4}$ $3{\times}10^{-4}$ $3{\times}10^{-4}$
Batch size 1024 1024 1024
Replay buffer size 500 k 500 k
Discount factor $\gamma$ 0.99 0.99 0.99
Polyak coefficient $\tau$ 0.005 0.005
Train freq / grad steps on-policy (256 steps) 4 / 6 4 / 8
Exploration term entropy $0.001$ $\alpha$ (auto) noise $\sigma=0.4$
Total env. steps $8{\times}10^{6}$ (4 workers)
Table 3: Hyper-parameters for the Dog/Car-2D environment.
Hyper-parameter PPO SAC TD3
Policy network [64, 64]6464[64,\,64][ 64 , 64 ] [64, 64]6464[64,\,64][ 64 , 64 ] [64, 64]6464[64,\,64][ 64 , 64 ]
Value / $Q$ network [64, 64]6464[64,\,64][ 64 , 64 ] [64, 64]6464[64,\,64][ 64 , 64 ] [64, 64]6464[64,\,64][ 64 , 64 ]
Learning rate $3{\times}10^{-4}$ $3{\times}10^{-4}$ $3{\times}10^{-4}$
Batch size 1024 1024 1024
Replay buffer size 500 k 500 k
Discount factor $\gamma$ 0.99 0.99 0.99
Polyak coefficient $\tau$ 0.005 0.005
Train freq / grad steps on-policy (256 steps) 4 / 6 4 / 8
Exploration term entropy $0.01$ $\alpha$ (auto) noise $\sigma=0.4$
Total env. steps $4{\times}10^{6}$ (4 workers)